OpenStudy (anonymous):

find the distance between p1(7,1/2pi) and p2(1,7/2pi) on the polar plane

OpenStudy (anonymous):

a.11 b.7 c.8 d.10

OpenStudy (michaelbp):

In order to solve this you have to use the law of cosines: c^2 = a^2 + b^2 - 2ab (cosC) The r values (7 and 1) are used for a and b in the equation and to find angle C you subtract the degrees. First though you have to convert radians into degrees which you get 90 degrees for 1/2 pie and 7/2 pie is 630 degrees or 270 degrees ( 7(180) /2 ). So the angle C is 270-90 = 180. Then you just plug in the values in the equation: c^2 = 7^2 + 1^2 - 2(7)(1) ( cos(180) ) c^2 = 49 + 1 - 14(-1) c^2 = sqrt(64) C= 8 so your answer is C

OpenStudy (anonymous):

wow! that was fast! thanks so much!

OpenStudy (michaelbp):

yw medal and fan please

OpenStudy (michaelbp):

thanks

OpenStudy (anonymous):

no prob

OpenStudy (anonymous):

could i ask you one more question